OpenStudy (mamamia922):

What is the range of values for the measurement 500±4%? I'm super confused on this and I would really appreciate any help!!!

jhonyy9 (jhonyy9):

than 500 mean the 100% how many will be 4% ? use the rule of 3 simple 500 ------ 100% x -------- 4% ------------------- x = 500*4/100 = ?

OpenStudy (mamamia922):

2000/100 which would = 20. Is that it??